OpenStudy (anonymous):

Using complete sentences, explain how to graph y = -1/4x + 8 using the slope-intercept method.

Parth (parthkohli):

\( \color{Black}{\begin{array} x x &y \\ 1 & ? \\ 2 &? \\ 3&? \end{array}}\) Make a table of values. Then mark the values on the graph, and draw a straight line though that.

Parth (parthkohli):

Another way: y intercept is 8(y = mx + b and b is the y intercept), and we know that when x is 0, then y is 8.

Parth (parthkohli):

As slope is -1/4, everytime we increase -1 in y, x will increase by 4. Find the different values.

OpenStudy (anonymous):

But how do explain this to a teacher with the current expression?

OpenStudy (anonymous):

Thats what I want to know.

Parth (parthkohli):

I just did :)

OpenStudy (anonymous):

y intercept is 8(y = mx + b and b is the y intercept), and we know that when x is 0, then y is 8.

OpenStudy (anonymous):

so thats the answer?